OpenStudy (anonymous):

A culture started with 4,000 bacteria. After 8 hours, it grew to 4,400 bacteria. Predict how many bacteria will be present after 9 hours. Round your answer to the nearest whole number. P=Ae^kt

OpenStudy (anonymous):

snap way \[\frac{4400}{4000}=1.1\] model as \[4000\times (1.1)^{\frac{t}{8}}\] at \(t=9\) you get \[4000\times (1.1)^{\frac{9}{8}}\] or if you prefer \[4400\times (1.1)^{\frac{1}{8}}\]

OpenStudy (anonymous):

otherwise you can do it the annoying way and solve \[4000e^{8k}=4400\] for \(k\) then write \[4000e^{9k}=...\]

OpenStudy (anonymous):

So the answer would be 4400 *1.1^1/3?

OpenStudy (anonymous):

your answer will be less accurate the second way because inevitable you will truncate \[k=\frac{\ln(1.1)}{8}\]

OpenStudy (anonymous):

no not \[4400(1.1)^{\frac{1}{3}}\] but rather \[4400(1.1)^{\frac{1}{8}}\]

OpenStudy (anonymous):

here it is done the snap way http://www.wolframalpha.com/input/?i=4400%281.1%29^%281%2F8%29

OpenStudy (anonymous):

pita way is to solve \[4000e^{8k}=4400\] for \(k\) via \[e^{8k}=1.1\] \[8k=\ln(1.1)\] \[k=\frac{\ln(1.1)}{8}=.0119\] \[P(t)=4000e^{.0119t}\] and then compute \[P(9)=4000e^{0.119\times 9}\] http://www.wolframalpha.com/input/?i=4000e^ {.0119*9%29 notice how close they are